# Temporal Interest Network for User Response Prediction

### v2

@@ Abstract.
User response prediction is essential in industrial recommendation systems, such as online display advertising. Among all the features in recommendation models, user behaviors are among the most critical. Many works have revealed that a user’s behavior reflects her interest in the candidate item, owing to the semantic or temporal correlation between behaviors and the candidate. While the literature has individually examined each of these correlations, researchers have yet to analyze them in combination, that is, the semantic-temporal correlation. We empirically measure this correlation and observe intuitive yet robust patterns. We then examine several popular user interest models and find that, surprisingly, none of them learn such correlation well.
@@
journalyear: 2024copyright: rightsretainedconference: Companion Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singapore.booktitle: Companion Proceedings of the ACM Web Conference 2024 (WWW ’24 Companion), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589335.3648340isbn: 979-8-4007-0172-6/24/05**footnotetext: Equal contribution
@@ 1. Introduction
Recommendation systems usually use categorical ID features to represent users, items, and the context, such as user demographic features, item attribute features, etc. Among them, users’ history behaviors are considered a crucial feature since they represent a user’s preference over items in the past and, therefore, may reflect her interest in the candidate item. Many works tried to capture such user interest by recurrent neural networks (Zhou et al., 2019; Hidasi et al., 2015), self-attention (Kang and McAuley, 2018; Sun et al., 2019; Chen et al., 2019), or target attention (Zhou et al., 2018, 2019; Feng et al., 2019; Chen et al., 2019). For instance, DIN (Zhou et al., 2018) models the semantic correlation between each behavior and the target by a target attention mechanism. SASRec (Kang and McAuley, 2018) captures the correlation between behaviors by self-attention. GRU4Rec (Hidasi et al., 2015) tries to capture the user’s interest within a session via recurrent neural networks (RNN).
@@
In particular, Each of the three components in TIN plays a critical role in acquiring the semantic-temporal correlation. The absence of target-aware temporal encoding renders the model incapable of capturing the target-relative temporal correlation. Without either target-aware attention or target-aware representation, TIN deteriorates to a 3rd-order or 2nd-order interaction model, thus unable to capture the quadruple semantic-temporal correlation.
To empirically validate the significance of each component, we conduct an ablation study on the Amazon dataset. The ablation of each component results in a notable decrease in the GAUC by 0.99%, 0.73%, and 0.61%, respectively. It is noteworthy that all existing methods can be considered as ablated variants of TIN, as they lack one or more of these critical components. The contribution of this paper can be summarized as follows:
@@
  • We conduct comprehensive experiments on two publicly available datasets, and the results demonstrate that TIN outperforms state-of-the-art user interest methods. In online A/B testing, TIN achieves 1.65% cost lift and 1.93% GMV lift over the base model, and it has been successfully deployed in Tencent’s online advertising platform.
@@ 2.1. Measurement of Ground Truth Semantic-Temporal Correlation
CTC is defined as the mutual information (Pan et al., 2018; Shwartz-Ziv and Tishby, 2017) between behaviors with category $c_{i}$ whilst occurring at position $p$: $\mathcal{X}_{C(X)=c_{i}\land P(X)=p}$ and the user response label ***The Amazon dataset only contains reviews, i.e., positive responses, on the target item. We generate negative responses by replacing the target item with a random item. on the target item with category $c_{t}$: $\mathcal{Y}_{C(Y)=c_{t}}$. Formally,
@@ 2.2. How do Existing Methods Capture Semantic-Temporal Pattern?
We are intrigued by how existing user interest methods capture the observed patterns. To investigate this, we select three popular user interest models: DIN (Zhou et al., 2018), SASRec (Kang and McAuley, 2018), and BST (Chen et al., 2019). Similarly, we evaluate the learned correlation of these models on the target category 674, as depicted in Fig. 1(b), Fig. 1(c) and Fig. 1(d).
@@ 3.2. Temporal Interest Module
where $\tilde{\bm{e}}_{i}=\bm{e}_{i}\oplus\bm{p}_{f(X_{i})}$ denotes the representation of the $i$-th behavior $X_{i}$, consisting of both semantic encoding $\bm{e}_{i}$ and target-aware temporal encoding $\bm{p}_{f(i)}$, similarly, $\tilde{\bm{v}}_{t}=\bm{v}_{i}\oplus\bm{p}_{f(X_{t})}$ denotes the representation of target item. $\alpha(\tilde{\bm{e}}_{i},\tilde{\bm{v}}_{t})$ and $(\tilde{\bm{e}}_{i}\odot\tilde{\bm{v}}_{t})$ denotes the target-aware attention and target-aware representation between the $i$-th behavior and the target, $\alpha(\cdot)$ denotes the softmax function, and $\odot$ denotes element-wise multiplication. We illustrated the architecture of TIM in Fig. 2(b).
@@ 3.3. Connection to Existing Methods
For instance, the code for TIN is ✓✓✓, indicating the presence of all three components. On the other hand, DIN’s code is ✗✓✓, as it lacks TI. Models such as GRU4Rec, SASRec, and BERT4Rec share the same code, ✓✗✓, as they incorporate TI and TR but lack TA. Conversely, other target attention methods like DIEN, DSIN, and BST have the code ✓✓✗, due to the failure of Concat & MLP to learn dot product (Rendle et al., 2020). Please refer to Table 5 in the Appendix for a comprehensive overview of each method’s code.
3.4. Connections between Temporal Interest Module and Scaled Dot-product Attention
There are several commonalities between the Temporal Interest Module (TIM) and Scaled Dot-Product Attention proposed in Transformer (Vaswani et al., 2017): 1. they both employ Scaled Dot-Product as the attention function; 2. they consider both semantic embedding as well as temporal embedding to represent each token or item. Using the Transformer terminology, $\tilde{\bm{v}}_{t}$ and $\tilde{\bm{e}}_{i}$ in target-aware attention corresponds to query (target query) and key (behavior key) in Dot-Product Attention, while $\tilde{\bm{e}}_{i}$ and $\tilde{\bm{v}}_{t}$ in target-aware representation corresponds to values (behavior value and target value, respectively).
@@ 4.1. RQ1: Overall Performance Evaluation
We compare TIN with the following user interest methods as baselines: Avg Pooling & Concat (Zhou et al., 2018), Avg Pooling & Product (Zhou et al., 2018), DIN’***the version described in the paper, consisting of only target-aware attention, DIN***the version implemented in the released code, which multiplies the attention-weighted pooling of behaviors with the target outside the attention, therefore consists of both target-aware attention and target-aware representation, GRU4Rec (Hidasi et al., 2015), SASRec, BERT4Rec (Sun et al., 2019), DIEN (Zhou et al., 2019), DSIN (Feng et al., 2019) and BST. Refer to the appendix for the experimental setting.
@@
All the remaining methods, namely GRU4Rec, SASRec, BERT4Rec DIEN, DSIN, and BST, take into account temporal information. It is worth noting that many of these methods outperform DIN on the Amazon dataset due to the presence of a strong target-aware temporal correlation, as shown in Fig. 1(a). However, on the Alibaba dataset, some of these methods fail to surpass the performance of DIN. This discrepancy can potentially be attributed to the dense temporal nature of behaviors in the Alibaba dataset, as discussed in detail in Section 4.4. The high density of behaviors in relation to time renders the position information less influential or useful in this context.
@@ 4.2.3. Disabling Target-aware Representation
4.3. RQ3: Measurement of Learned Semantic-Temporal Correlation
@@
where $z$ represents the attention logits and $\bm{r}$ denotes the representation embedding. The exponential of the logits $e^{z}$ is taken because the softmax activation function normalizes the exponential of $z$ across all behaviors. The $L_{2}$ norm of the representation, $\|\bm{r}\|$, is chosen to measure its strength, although other norms could also be considered. For instance, the learned semantic-temporal correlation of TIN can be defined as:
@@
4.4. RQ4: Comparison of Temporal Encoding Methods
To evaluate the effectiveness of these two temporal encoding methods, we partitioned the Amazon dataset based on the sequence length of samples and conducted a comparative analysis. As presented in Table 3, TTE-P consistently outperforms COE across all sequence length ranges, with the performance gap widening for longer sequences. Unfortunately, we are unable to perform a similar comparison on the Alibaba dataset since most sequences are padded to a fixed length of 50.
Table 3. Performance evaluation of various temporal encoding methods on the Amazon dataset with various behavior sequence length ranges.
@@ 4.4. RQ4: Comparison of Temporal Encoding Methods
Next, we present the comparison results of TTE-P, TTE-T, and Timestamp Encoding on both datasets in Table 4. To discretize time, we define the function $b(\cdot)$ as an equal frequency binning approach, ensuring that each range contains an equal number of samples. In this study, we split the time into 10 intervals. Upon replacing TTE-P with TTE-T, we observe an increase in the GAUC by 8.6e-3 on the Amazon dataset and a slight drop of 0.8e-3 on the Alibaba dataset. This discrepancy can be attributed to the distinct time ranges of behaviors in the two datasets. The Amazon dataset encompasses review data spanning from 1996 to 2014, whereas the Alibaba dataset comprises user behaviors recorded over a continuous period of 8 days. The extensive time range of the Amazon dataset renders TTE-T more informative than TTE-P. Conversely, on the Alibaba dataset with a relatively short time range, TTE-P proves to be more effective in capturing temporal patterns.
Table 4. Performance evaluation of various encoding methods on two datasets.
@@
4.5. Summary of models regarding the three components
Table 5 presents a comprehensive summary of existing models, including TIN, with respect to Temporal Information, Target-aware Attention, and Target-aware Representation.
DIN consists solely of Target-aware Attention (TA), while DIN’ incorporates both TA and Target-aware Representation (TR). In DIEN (Zhou et al., 2019), the Interest Extractor Layer (ExtLayer) and Interest Evolution Layer (EvoLayer) capture Temporal Information (TI) using the Gated Recurrent Unit (GRU). However, ExtLayer lacks both TA and TR, while EvoLayer lacks TR. DSIN (Feng et al., 2019) employs the Session Interest Extractor (SIE), which applies self-attention over historical behaviors. The Session Interest Activating (SIA) layer computes an attention weight for each behavior based on the behavior itself and the target item. DSIN then feeds the output of SIA to a Concatenation and Multi-Layer Perceptron (MLP) layer. Consequently, DSIN incorporates TI through self-attention and TA via SIA, but it lacks TR due to the usage of Concatenation and MLP. BST (Chen et al., 2019) combines both TI (achieved through position encoding) and TA using the Transformer model, which can be viewed as an extension of the DIN model. However, BST simply concatenates the output of the Transformer, which represents an attentive pooling of behavior embeddings, with the target item embedding. This concatenated representation then undergoes several MLP layers to make predictions. As mentioned earlier, the Concatenation and MLP are hard to learn the dot product (interaction) effectively (Rendle et al., 2020).
@@ 4.6. Online A/B Testing
We deploy the following techniques to meet the online real-time inference latency and throughput criterion for real-world large-scale recommendation systems: a) Mixed-precision Inference: Our model employs the in-house inference engine HCF for online inference, supporting INT8, FP16, and FP32 mixed-precision inference. b) Automatic Feature Selection: We have transitioned from traditional greedy methods for feature selection (Deng et al., 2021) to feature-set selection (Wei et al., 2023), which significantly reduces the number of features. c) Automatic Neural Network Architecture Search: Through AutoDL, we select the optimal feature cross sub-network (Wei et al., 2024), reducing unnecessary feature combinations and improving performance.
@@ 5. Related Work
Another line of research is sequential recommenders (Tang and Wang, 2018; Kang and McAuley, 2018; Yuan et al., 2019; Li et al., 2022; Yuan et al., 2022; Zhang et al., 2022; Su et al., 2023a; Yuan et al., 2022; Hu et al., 2022; Liu et al., 2021b; Xie et al., 2022c; Zhou et al., 2020; Xie et al., 2022b; Lin et al., 2024), which aim to capture temporal correlations using either Recurrent Neural Networks (RNN) (Hidasi et al., 2015; Wu et al., 2017; Yu et al., 2016) or Self-Attention (Kang and McAuley, 2018; Sun et al., 2019; Liu et al., 2021a; Wu et al., 2020). However, these models lack target-aware attention, limiting them to only being able to capture 2nd or 3rd-order interactions and, therefore, unable to learn quadruple semantic-temporal correlation.
@@ 6. Conclusion
This paper investigates the crucial semantic-temporal correlation between user behaviors and the target. Examination of existing methods reveals that they fail to capture such correlation. We propose the Temporal Interest Network (TIN), which incorporates target-aware temporal encoding, target-aware attention, and target-aware representation. Comprehensive experiments on two public datasets demonstrate the superiority of TIN over the best-performing baselines. TIN has been successfully deployed to WeChat Moments on Tencent’s advertising platform.
@@ A.1. Experimental Setting
The Amazon***http://jmcauley.ucsd.edu/data/amazon/ dataset is derived from real product reviews and ratings on the Amazon e-commerce website. We conduct experiments on the Electronics subset, which contains 192,403 users, 63,001 goods, 801 categories and 1,689,188 samples. Each user in the dataset has a minimum of 5 reviews with various items. We adopt the widely employed leave-one-out strategy (Zhou et al., 2018), that is, we predict the user’s feedback for the $n$-th item based on the user’s first $n-1$ interaction records. The negative samples are generated by replacing the target item in positive samples with a random item from the entire item collection.
The Alibaba***https://tianchi.aliyun.com/dataset/56 dataset encompasses a comprehensive collection of data obtained from a random selection of 1.14 million users on Taobao’s website, resulting in an 8-day advertising display /click log comprising 26 million records. We partition the dataset as follows: the samples from the first six days (2017-05-06 to 2017-05-11) are allocated for training purposes, the samples occurring on 2017-05-12 are utilized for validation, and the samples from the final day (2017-05-13) are reserved for testing. We follow the data preprocessing procedure of DSIN (Feng et al., 2019).
@@
For both datasets, we set the embedding dimension $d$ to 64 for all feature embeddings. The hidden layers of the two-layer MLP have dimensions of 80 and 40 for the Amazon dataset, and 200 and 80 for the Alibaba dataset. As the optimizer, we employ Adam (Kingma and Ba, 2014) with a learning rate of 0.001 for the Amazon dataset, and Adagrad (Duchi et al., 2011) with a learning rate of 0.01 for the Alibaba dataset. To ensure consistency and comparability, we maintain the remaining settings identical to those used in DIN on the Amazon dataset, and DSIN***https://github.com/shenweichen/DSIN/tree/master/code on the Alibaba dataset. We evaluate the performance using Logloss and GAUC. To account for variability, each experiment is repeated five times, and we report the mean and standard deviation of the results.

### v3

@@ Abstract.
User behaviors are among the most critical features for user response prediction in recommendation systems. Many works have revealed that a user’s behavior reflects her interest in the candidate item, owing to the semantic or temporal correlation between them. While the literature has individually examined each of these correlations, researchers have yet to analyze them in combination, that is, the semantic-temporal correlation. We empirically measure this correlation and observe intuitive yet robust patterns. We then examine several popular user interest models and find that, surprisingly, none of them learn such correlation well.
@@
journalyear: 2024copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Companion Proceedings of the ACM Web Conference 2024 (WWW ’24 Companion), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589335.3648340isbn: 979-8-4007-0172-6/24/05ccs: Computing methodologies Semantic networks
@@ 1. Introduction
Recommendation systems usually use categorical ID features to represent users, items, and the context, such as user demographic features, item attribute features, etc. Among them, users’ history behaviors are crucial since they represent a user’s preference over items in the past and, therefore, may reflect her interest in the candidate item. Many works tried to capture such user interest by recurrent neural networks (Zhou et al., 2019; Hidasi et al., 2015), self-attention (Kang and McAuley, 2018; Sun et al., 2019; Chen et al., 2019), or target attention (Zhou et al., 2018, 2019; Feng et al., 2019; Chen et al., 2019).
@@
Each of the three components in TIN is critical in acquiring the semantic-temporal correlation. To empirically validate the significance of each component, we conduct an ablation study on the Amazon dataset. The ablation of each component results in a notable decrease in the GAUC by 0.99%, 0.73%, and 0.61%, respectively. It is noteworthy that all existing methods can be considered as ablated variants of TIN, as they lack one or more of these critical components. The contribution of this paper can be summarized as follows:
@@
  • We conduct comprehensive experiments on two publicly available datasets, and the results demonstrate that TIN outperforms state-of-the-art user interest methods.
@@ 2.1. Measurement of Ground Truth Semantic-Temporal Correlation
CTC is defined as the mutual information (Pan et al., 2018; Shwartz-Ziv and Tishby, 2017) between behaviors with category $c_{i}$ whilst occurring at position $p$: $\mathcal{X}_{C(X)=c_{i}\land P(X)=p}$ and the user response label  [^1] The Amazon dataset only contains reviews, i.e., positive responses, on the target item. We generate negative responses by replacing the target item with a random item. on the target item with category $c_{t}$: $\mathcal{Y}_{C(Y)=c_{t}}$. Formally,
@@ 2.2. How do Existing Methods Capture Semantic-Temporal Pattern?
We are intrigued by how existing user interest methods capture the observed patterns. To investigate this, we select three popular user interest models: DIN (Zhou et al., 2018), SASRec (Kang and McAuley, 2018), and BST (Chen et al., 2019), and evaluate the learned correlation of these models on the target category 674, as depicted in Fig. 1(b), Fig. 1(c) and Fig. 1(d).
@@ 3.1. Temporal Interest Module
where $\tilde{\bm{e}}_{i}=\bm{e}_{i}\oplus\bm{p}_{f(X_{i})}$ denotes the representation of the $i$-th behavior $X_{i}$, consisting of both semantic encoding $\bm{e}_{i}$ and target-aware temporal encoding $\bm{p}_{f(i)}$, similarly, $\tilde{\bm{v}}_{t}=\bm{v}_{t}\oplus\bm{p}_{f(X_{t})}$ denotes the representation of target item. $\alpha(\tilde{\bm{e}}_{i},\tilde{\bm{v}}_{t})$ and $(\tilde{\bm{e}}_{i}\odot\tilde{\bm{v}}_{t})$ denotes the target-aware attention and target-aware representation between the $i$-th behavior and the target, $\alpha(\cdot)$ denotes the softmax function, and $\odot$ denotes element-wise multiplication. We illustrated the architecture of TIM in Fig. 2(b).
@@ 3.2. Connection to Existing Methods
For instance, the code for TIN is ✓✓✓, indicating the presence of all three components. On the other hand, DIN’s code is ✗✓✓, as it lacks TI. Models such as GRU4Rec, SASRec, and BERT4Rec share the same code, ✓✗✓, as they incorporate TI and TR but lack TA. Conversely, other target attention methods like DIEN, DSIN, and BST have the code ✓✓✗, due to the failure of Concat & MLP to learn dot product (Rendle et al., 2020). Please refer to Table 5 for a comprehensive overview of each method’s code.
3.3. Connections to Transformer
There are several commonalities between the Temporal Interest Module (TIM) and Scaled Dot-Product Attention proposed in Transformer (Vaswani et al., 2017): 1. they both employ Scaled Dot-Product as the attention function; 2. they consider both semantic embedding as well as temporal embedding to represent each token or item. Using the Transformer terminology, $\tilde{\bm{v}}_{t}$ and $\tilde{\bm{e}}_{i}$ in target-aware attention corresponds to query (target query) and key (behavior key) in Dot-Product Attention, while $\tilde{\bm{e}}_{i}$ and $\tilde{\bm{v}}_{t}$ in target-aware representation corresponds to values (behavior value and target value).
@@ 4.1. RQ1: Overall Performance Evaluation
We compare TIN with the following user interest methods as baselines: Avg Pooling & Concat (Zhou et al., 2018), Avg Pooling & Product (Zhou et al., 2018), DIN’ [^1] DIN’ denotes the architecture described in the paper, consisting of only target-aware attention, DIN [^2] DIN denotes the architecture implemented in the released code, which multiplies the attention-weighted pooling of behaviors with the target outside the attention, therefore consists of both target-aware attention and target-aware representation, GRU4Rec (Hidasi et al., 2015), SASRec, BERT4Rec (Sun et al., 2019), DIEN (Zhou et al., 2019), DSIN (Feng et al., 2019) and BST. Refer to the appendix for the experimental setting.
@@
All the remaining methods, namely GRU4Rec, SASRec, BERT4Rec DIEN, DSIN, and BST, take into account temporal information. It is worth noting that many of these methods outperform DIN on the Amazon dataset due to the presence of a strong target-aware temporal correlation, as shown in Fig. 1(a). However, on the Alibaba dataset, some of these methods fail to surpass DIN. This discrepancy can potentially be attributed to the dense temporal nature of behaviors in the Alibaba dataset, as discussed in detail in Section 4.4. The high density of behaviors in relation to time renders the position information less influential or useful in this context.
@@ 4.2.3. Disabling Target-aware Representation
4.3. RQ3: Measurement of Learned Correlation
@@
where $z$ represents the attention logits and $\bm{r}$ denotes the representation embedding. We chose this form since there are connections between the mutual information and the parameterized (intermediate) output of the interaction of two variables (Oord et al., 2018). In our model, the intermediate parameterized output of the interaction before the MLPs is proportional to $e^{z}\cdot\bm{r}$. For instance, the learned semantic-temporal correlation of TIN can be defined as:
@@
Table 3. Performance evaluation of temporal encoding methods on the Amazon dataset with various behavior sequence length ranges.
@@
Table 4. Performance evaluation of various encoding methods.
@@
4.4. RQ4: Temporal Encoding Methods
To evaluate these two temporal encoding methods, we partitioned the Amazon dataset based on the sequence length and conducted a comparative analysis. As shown in Table 3, TTE-P consistently outperforms COE across all sequence length ranges, with the performance gap widening for longer sequences. Unfortunately, we are unable to perform a similar comparison on the Alibaba dataset since most sequences are padded to a fixed length of 50.
Next, we present the comparison results of TTE-P, TTE-T, and Timestamp Encoding on both datasets in Table 4. Specifically, TTE-T split the time interval into 10 buckets with equal frequency binning function. Upon replacing TTE-P with TTE-T, we observe an increase in the GAUC by 8.6e-3 on the Amazon dataset and a slight drop of 0.8e-3 on the Alibaba dataset. This discrepancy can be attributed to the distinct time ranges of behaviors in the two datasets. The Amazon dataset encompasses review data from 1996 to 2014, whereas the Alibaba dataset comprises user behaviors recorded over a continuous period of 8 days. The larger time range of the Amazon dataset renders TTE-T more informative than TTE-P. Conversely, on the Alibaba dataset with a relatively short time range, TTE-P proves to be more effective in capturing temporal patterns.
4.5. Existing models regarding the three components
Table 5 presents a comprehensive summary of existing models, including TIN, with respect to Temporal Information (TI), Target-aware Attention (TA), and Target-aware Representation (TR). DIN consists solely of TA, while DIN’ incorporates both TA and TR. In DIEN (Zhou et al., 2019), the Interest Extractor Layer (ExtLayer) and Interest Evolution Layer (EvoLayer) capture TI using the Gated Recurrent Unit (GRU). However, ExtLayer lacks both TA and TR, while EvoLayer lacks TR. DSIN (Feng et al., 2019) employs the Session Interest Extractor (SIE), which applies self-attention over historical behaviors. The Session Interest Activating (SIA) layer computes an attention weight for each behavior based on the behavior itself and the target item. DSIN then feeds the output of SIA to a Concatenation and Multi-Layer Perceptron (MLP) layer. Consequently, DSIN incorporates TI through self-attention and TA via SIA, but it lacks TR due to the usage of Concatenation and MLP. BST (Chen et al., 2019) combines both TI (achieved through position encoding) and TA using the Transformer model, which can be viewed as an extension of the DIN model. However, BST simply concatenates the output of the Transformer, which represents an attentive pooling of behavior embeddings, with the target item embedding. This concatenated representation then undergoes several MLP layers to make predictions. As mentioned earlier, the Concatenation and MLP are hard to learn the dot product (interaction) effectively (Rendle et al., 2020).
@@ 5. Related Work
Another line of research is sequential recommenders (Tang and Wang, 2018; Kang and McAuley, 2018; Yuan et al., 2019; Su et al., 2023a; Hu et al., 2022; Liu et al., 2021b; Xie et al., 2022c; Zhou et al., 2020; Xie et al., 2022b; Lin et al., 2024), which aim to capture temporal correlations using either Recurrent Neural Networks (RNN) (Hidasi et al., 2015; Wu et al., 2017; Yu et al., 2016) or Self-Attention (Kang and McAuley, 2018; Sun et al., 2019; Liu et al., 2021a; Wu et al., 2020). However, these models lack target-aware attention, limiting them to only being able to capture 2nd or 3rd-order interactions and, therefore, unable to learn quadruple semantic-temporal correlation.
@@ 6. Conclusion
This paper investigates the crucial semantic-temporal correlation between user behaviors and the target. Examination of existing methods reveals that they fail to capture such correlation. We propose the Temporal Interest Network (TIN) to capture such correlation, which incorporates target-aware temporal encoding, target-aware attention, and target-aware representation. Comprehensive experiments on two public datasets demonstrate the superiority of TIN over the best-performing baselines. TIN has been successfully deployed to WeChat Moments on Tencent’s advertising platform.
@@ 7. Acknowledgement
This work was supported by the National Natural Science Foundation of China [U23A20309, 62272302, 62172276, 62372296], Shanghai Municipal Science and Technology Major Project [2021SHZDZX-0102], and the Tencent Rhinoceros Project. Haolin Zhou, Xinyi Zhou, Xiaofeng Gao and Guihai Chen are in the MoE Key Lab of Artificial Intelligence, Department of Computer Science and Engineering, Shanghai Jiao Tong University, Shanghai, China. Xiaofeng Gao is the corresponding author.
@@ A.1. Experimental Setting
The Amazon [^2] http://jmcauley.ucsd.edu/data/amazon/ dataset is derived from real product reviews and ratings on the Amazon e-commerce website. We conduct experiments on the Electronics subset, which contains 192,403 users, 63,001 goods, 801 categories and 1,689,188 samples. Each user in the dataset has a minimum of 5 reviews with various items. We adopt the widely employed leave-one-out strategy (Zhou et al., 2018), that is, we predict the user’s feedback for the $n$-th item based on the user’s first $n-1$ interaction records. The negative samples are generated by replacing the target item in positive samples with a random item from the entire item collection.
The Alibaba [^3] https://tianchi.aliyun.com/dataset/56 dataset encompasses a comprehensive collection of data obtained from a random selection of 1.14 million users on Taobao’s website, resulting in an 8-day advertising display /click log comprising 26 million records. We partition the dataset as follows: the samples from the first six days (2017-05-06 to 2017-05-11) are allocated for training purposes, the samples occurring on 2017-05-12 are utilized for validation, and the samples from the final day (2017-05-13) are reserved for testing. We follow the data preprocessing procedure of DSIN (Feng et al., 2019).
@@
For both datasets, we set the embedding dimension $d$ to 64 for all feature embeddings. The hidden layers of the two-layer MLP have dimensions of 80 and 40 for the Amazon dataset, and 200 and 80 for the Alibaba dataset. As the optimizer, we employ Adam (Kingma and Ba, 2014) with a learning rate of 0.001 for the Amazon dataset, and Adagrad (Duchi et al., 2011) with a learning rate of 0.01 for the Alibaba dataset. To ensure consistency and comparability, we maintain the remaining settings identical to those used in DIN on the Amazon dataset, and DSIN [^4] https://github.com/shenweichen/DSIN/tree/master/code on the Alibaba dataset. We evaluate the performance using Logloss and GAUC. To account for variability, each experiment is repeated five times, and we report the mean and standard deviation of the results.